OpenStudy (anonymous):

Calc derivatives would f(x) = e^[(4x-1)^2] just be f'(x) = e^[(4x-1)^2] ?

OpenStudy (kirbykirby):

No You need to apply the chain rule

OpenStudy (anonymous):

let me work this out for a second

OpenStudy (kirbykirby):

ok

OpenStudy (anonymous):

wait, I am lost :(

OpenStudy (anonymous):

I would start by doing (4x-1)^(2) (e) ?

OpenStudy (anonymous):

i mean (4x-1)^(2) (e)^[(4x-1)^(2) -1]

OpenStudy (kirbykirby):

\[\frac{d}{du}e^u = e^u \cdot u' \]

OpenStudy (anonymous):

no its not like that,use this rule \[\huge (e^{f(x)})'=f'(x)e^{f(x)}\]

OpenStudy (anonymous):

so u need to derive \[(4x-1)^2\]

OpenStudy (anonymous):

ok, 2(4x-1)

OpenStudy (anonymous):

use chain rule

OpenStudy (kirbykirby):

\[\frac{d}{dx} e^{(4x-1)^2}=e^{(4x-1)^2}\cdot \frac{d}{dx}(4x-1)^2\]

OpenStudy (kirbykirby):

Don't forget to find the derivative inside of (4x-1) when using the chain rule

OpenStudy (anonymous):

Oh, thanks. I almost forgot

OpenStudy (anonymous):

Ok, so would the answer be \[e ^{4x-1^{2}} (8x+2)\]

OpenStudy (kirbykirby):

Euh not quite

OpenStudy (anonymous):

What did i miss? Is it just simplification?

OpenStudy (anonymous):

where did + come from

OpenStudy (kirbykirby):

the (8x+2) factor is wrong. And you forgot the parentheses in the exponential (4x-1)^2

OpenStudy (anonymous):

so we shud have \[(2)(4x-1)(4)=8(4x-1)\]

OpenStudy (kirbykirby):

^yup

OpenStudy (anonymous):

Oh, i see it. I skipped a step

OpenStudy (anonymous):

Wait, why is it (2)(4x−1)(4) and not (2)(4x−1) + (4) ?

OpenStudy (anonymous):

never mind, I see. I mixed up my rules

OpenStudy (anonymous):

Thank you both. I don't know who to give the medal to, so I will give it to the lower rank to help them out, but thank you both!!!

OpenStudy (kirbykirby):

I'll give a medal too haha